OpenStudy (anonymous):

I am doing Coursera and there is an Interger Exponents question I am confused at> Expression (4x^-3y^5z^-2)^-2/6y^-5z^5x^-8 equals Kz^r/x^sy^t where K= ??> What do they mean by quote "where K =?

OpenStudy (whpalmer4):

\[\frac{(4x^{-3}y^5z^{-2})^{-2}}{6y^{-5}z^5x^{-8}} = \frac{Kz^2}{x^sy^t}\] After you simplify the thing on the left, you'll be left with something like what is on the right, except that K will be an expression, and they want you to find it.